OpenStudy (calculusxy):

evaluate this step by step:

OpenStudy (calculusxy):

\[5^{2}\div(-5)(-3)-2(4\times3)\]

OpenStudy (calculusxy):

@shamil98

OpenStudy (calculusxy):

@ganeshie8

OpenStudy (shamil98):

25 divided by 15- 2(12)

OpenStudy (shamil98):

25/15 - 24

OpenStudy (calculusxy):

25/15=1.6666... how does that do the work? @shamil98

OpenStudy (shamil98):

25/15 = 5/3 5/3 - 24 5/3 - 72/3 = -67/3

OpenStudy (calculusxy):

I have to write this in numbers, not fractions.

OpenStudy (shamil98):

PEMDAS, you evaluate what is in parentheses first. then do the problem from left to right..

OpenStudy (shamil98):

well actually im wrong

OpenStudy (shamil98):

25/-5 = -5 -5(-3) - 2(12) 15 - 24 = -9

OpenStudy (calculusxy):

how did u get -5 from?

OpenStudy (shamil98):

if it were ((-5)(-3)) it would've been -67/3 but it is not..

OpenStudy (shamil98):

5^2 = 25 25/-5 = -5

OpenStudy (calculusxy):

thx

OpenStudy (shamil98):

mhm